OpenStudy (anonymous):

help please solve the compound inequality 2x-5/4 <= - 6 or 2x-5/4 > 5

OpenStudy (anonymous):

help please solve the compound inequality 2x-5/4 <= - 6 or 2x-5/4 > 5

OpenStudy (anonymous):

Solve for x individually in each one, just as if it were an equal sign, and when you have isolated both of your x's you have solved it.

OpenStudy (anonymous):

i'm struggling with this could you show me how to work this please

OpenStudy (anonymous):

dfyodor could you show me how to work this please

OpenStudy (anonymous):

All right.

OpenStudy (anonymous):

If I gave you this, could you solve it? \[2x-\frac 5 4 = -6\]

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

In which case, I regret to say that you lack the prerequisites to solve the problem. This is not an appropriate place to teach basic prealgebra skills. You should consult with a teacher, or someone who can show you first-hand.

OpenStudy (anonymous):

this is the first part. \[2x-\frac{5}{4}<=-6\]\[8x−5<=−24\]\[8x<=−19\]\[x<=−\frac{19}{8}\]The second part is: \[2x−\frac{5}{4}>5\]\[8x−5>20\]\[8x>25\]\[x>\frac{25}{8}\]Now the last part \[\frac{−19}{8}>=x>\frac{25}{8}\] There is your answer.